# Preconditioning FEM discretisations of the high-frequency Maxwell equations by either perturbing the coefficients or adding absorption

### v1

@@ Abstract
We prove bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ where $\mathsf{A}_{\ell}$, $\ell=1,2$, are the Galerkin matrices corresponding to finite-element discretisations of the time-harmonic Maxwell equations $k^{-2}{\rm curl}(\mu_{\ell}^{-1}{\rm curl}E_{\ell})-\epsilon_{\ell}E_{\ell}=f$; i.e., we consider the situation where the Maxwell FEM matrix is preconditioned by the FEM matrix arising from the same Maxwell problem but with different coefficients. An important special case is when the perturbation consists of adding absorption (in the spirit of “shifted Laplacian preconditioning” for the Helmholtz equation). The results of this paper are the Maxwell analogues of the Helmholtz results in [16, 20], and confirm a conjecture in the recent preprint [26]. These results are obtained by putting the Maxwell problem in an abstract framework that also includes the Helmholtz problem; as a byproduct we weaken the assumptions required to obtain the Helmholtz results in [16, 20].
1 Context and motivation
The design of good preconditioners for the linear systems arising from finite-element discretisations of high-frequency time-harmonic wave problems – such as the Helmholtz or time-harmonic Maxwell equations – is a longstanding open problem; see, e.g., the reviews [9, 12, 18, 17] and the references therein. This paper considers the theory of the following two, related, questions involving preconditioning these linear systems.
@@ 1 Context and motivation
The results of [16] were generalised in [20] to bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$, where $\mathsf{A}_{\ell}$, $\ell=1,2$, are the Galerkin matrices of $k^{-2}\nabla\cdot(A_{\ell}\nabla u_{\ell})+n_{\ell}u_{\ell}=f$. The motivation for studying this situation comes from uncertainty quantification (UQ): calculating quantities of interest for the solution of the Helmholtz equation $k^{-2}\nabla\cdot(A\nabla u)+nu=f$ with random coefficients requires the solution of many deterministic Helmholtz problems, each with different coefficients $A$ and $n$. Bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ then indicate to what extent a previously-calculated inverse of one of the Galerkin matrices can be used as a preconditioner for other Galerkin matrices; see [32, §4.6] and the recent work [35] for UQ algorithms using this idea.
@@
Outline of the paper.
@@
§2 states the abstract assumptions under which the main results are proved, and §3 describes Helmholtz and Maxwell problems that are covered by these abstract assumptions. §4 recaps properties of the inf-sup constant of a sesquilinear form and its $k$-dependence for Helmholtz and Maxwell. §5 states the main results, and §6 proves them.
2 Abstract assumptions
${\cal H}\subset{\cal H}_{0}$ are Hilbert spaces with $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$. We identify ${\cal H}_{0}$ with its dual so that ${\cal H}\subset{\cal H}_{0}\subset{\cal H}^{*}$. The linear operator ${\mathcal{D}}:{\cal H}\to{\cal H}_{0}$ with $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$. $b(\cdot,\cdot)$ is a continuous sesquilinear form on ${\cal H}$,
@@ 2 Abstract assumptions
We make four remarks.
  • In all examples below, $\|v\|^{2}_{{\cal H}}$ equals $\|{\mathcal{D}}v\|_{{\cal H}_{0}}^{2}+\|v\|_{{\cal H}_{0}}^{2}$ plus possibly an additional term; thus the condition $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$ is immediately satisfied.
@@
3 Helmholtz and Maxwell problems satisfying the abstract assumptions
3.1 Helmholtz problems
In all the Helmholtz problems, $\Omega$ is a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (scalar valued), ${\mathcal{D}}:=k^{-1}\nabla$, and
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}\nabla v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}$
(so that immediately $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$). In addition $\mu_{\ell}^{-1}$ are bounded, symmetric, matrix functions on $\Omega$ with $\Re\mu_{\ell}^{-1}>0$ (in the sense of quadratic forms) and $\epsilon_{\ell}$ are bounded, scalar-valued functions on $\Omega$. Note that these properties of $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are satisfied by the coefficients coming from a radial PML by, e.g., [14, Lemma 2.3].
3.1.1 Helmholtz problem with Dirichlet boundary conditions on part of the boundary (including PML approximation to scattering by a Dirichlet or Neumann obstacle)
:=${\cal H}:=\big{\{}u\in H^{1}(\Omega)\,:\,\text{ $u=0$ on at least one connected component of $\partial\Omega$}\big{\}}$
and $b(\cdot,\cdot)=0$. In this case, the sesquilinear form corresponds to the Helmholtz operator
$-k^{-2}\nabla\cdot(\mu_{\ell}^{-1}\nabla)-\epsilon_{\ell},$ (3.1)
with zero Dirichlet boundary conditions on at least one connected component of $\partial\Omega$ and zero Neumann boundary conditions on the rest.
3.1.2 Helmholtz problem with impedance boundary conditions on part of the boundary
With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$, and $\Gamma_{\rm D}$ a (possibly empty) connected component of $\partial\Omega$ that is not equal to $\Gamma_{\rm imp}$,
:=${\cal H}:=\big{\{}v\in H^{1}(\Omega)\,:\,v=0\text{ on }\Gamma_{\rm D}\big{\}},$
@@ 3.1.2 Helmholtz problem with impedance boundary conditions on part of the boundary
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$. To see that (2.2) is satisfied, observe that $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3) (with the term on $\Gamma_{\rm imp}$ playing no role).
In this case, the sesquilinear form corresponds to the Helmholtz operator (3.1) with the impedance boundary condition
@@
$u=0$ on $\Gamma_{\rm D}$, and $k^{-1}\partial_{n,\mu^{-1}}u=g_{\rm N}$ on $\partial\Omega\setminus(\Gamma_{\rm D}\cup\Gamma_{\rm imp})$ (where the data $g_{\rm imp}$ and $g_{\rm N}$ depend on the given right-hand side). Here $\partial_{n,\mu^{-1}}$ is the conormal derivative; recall that this is such that $\partial_{n,\mu^{-1}}u:=n\cdot\mu^{-1}\nabla u$ when $u\in H^{2}(\Omega)$, where $n$ is the outward-pointing unit normal vector to $\partial\Omega$ (see, e.g., [28, Lemmas 4.2 and 4.3]).
3.1.3 Helmholtz problem with exact Dirichlet-to-Neumann map imposed on the boundary of a ball
$\Omega=\{x:|x|\leq R\}\setminus\overline{\Omega_{-}}$, where $\Omega_{-}$ is a bounded Lipschitz open set with connected open complement (so that the scattering problem with obstacle $\Omega_{-}$ makes sense). ${\cal H}$ equals either $H^{1}(\Omega)$ or $\{v\in H^{1}(\Omega)\,:\,v=0\text{ on }\partial\Omega_{-}\}$.
$b(u,v):=-{\rm i}k^{-1}\big{(}{\rm DtN}_{k}u,v\big{)}_{L^{2}(\Gamma_{\rm imp})},$
where ${\rm DtN}_{k}$ is the exact Dirichlet-to-Neumann map for the Helmholtz equation in the exterior of $\{x:|x|\leq R\}$; see, e.g., [1, Equations 3.5 and 3.6] for explicit expressions in terms of Fourier series (in 2-d) or spherical harmonics (in 3-d) and Bessel and Hankel functions. To see that (2.2) is satisfied, recall that $\Re(-{\rm i}{\rm DtN}_{k})\geq 0$ by, e.g., [1, Corollary 3.1], so that $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3).
3.2 Maxwell problems
In all the Maxwell problems $\Omega$ is a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (vector valued), ${\mathcal{D}}:=k^{-1}{\rm curl}$, and the sesquilinear form corresponds to the Maxwell operator
$k^{-2}{\rm curl}\big{(}\mu^{-1}_{\ell}{\rm curl}\big{)}-\epsilon_{\ell}.$ (3.2)
3.2.1 Maxwell problem with PEC boundary condition (including PML approximation to scattering by a PEC obstacle)
@@ 3.2 Maxwell problems
where $v_{T}$ is the tangential trace such that $v_{T}:=(n\times v)\times n$ for smooth $v$, where $n$ is the outward-pointing unit normal vector (see, e.g., [31, Equation 3.46 and Theorem 3.31]).
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}$
and $b(\cdot,\cdot)=0$. $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are bounded, symmetric, matrix functions on $\Omega$ with $\Re\mu_{\ell}^{-1}>0$ and $\Re\epsilon_{\ell}>0$ (in the sense of quadratic forms). (Note that these properties are satisfied by the coefficients coming from a radial PML by, e.g., [3, Appendix A].)
3.2.2 Maxwell problem with impedance boundary conditions on part of the boundary
With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$,
:=${\cal H}:=\big{\{}u\in H^{1}(\Omega)\,:\,\text{ $u_{T}=0$ on $\partial\Omega\setminus\Gamma_{\rm imp}$ and $u_{T}\in L^{2}(\Gamma_{\rm imp})$}\big{\}},$
$\left\|v\right\|^{2}_{{\cal H}}:=k^{-2}\left\|{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}+k^{-1}\left\|v_{T}\right\|^{2}_{L^{2}(\Gamma_{\rm imp})},$
@@ 3.2.2 Maxwell problem with impedance boundary conditions on part of the boundary
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$. $\mu_{\ell}^{-1}$ are real-valued bounded, symmetric, matrix functions on $\Omega$ with $\mu_{\ell}^{-1}>0$ (in the sense of quadratic forms). $\epsilon_{\ell}$ are real-valued, bounded, symmetric matrix functions on $\Omega$.
To see that (2.2) is satisfied, observe that, since $\mu^{-1}_{\ell}$ and $\epsilon_{\ell}$ are all real, $\Im(a(v,v))$ consists only of $\Im(b(v,v))$, which is bounded below by $k^{-1}({\rm essinf}_{\Gamma_{\rm imp}}\theta)\|v\|^{2}_{L^{2}(\Gamma_{\rm imp})}$.
In this case, the sesquilinear form corresponds to the Maxwell operator (3.2) with the impedance boundary condition
@@
Most papers on the numerical analysis of the Helmholtz and time-harmonic Maxwell equations write these equations as $(\Delta+k^{2})u=f$ and $({\rm curl}\,{\rm curl}-k^{2})E=f$ (in the constant coefficient case) and use the norms $\|v\|^{2}_{H^{1}}=\|\nabla v\|^{2}_{L^{2}}+k^{2}\|v\|^{2}_{L^{2}}$ and $\|w\|_{H({\rm curl})}^{2}=\|{\rm curl}w\|^{2}_{L^{2}}+k^{2}\|w\|^{2}_{L^{2}}$. The advantage of rescaling the PDEs and norms so that every derivative appears with a $k^{-1}$ is that none of the constants in §2 (i.e. $C_{\rm G1,\ell},C_{\rm G2,\ell},$ and $C_{\rm cont,\ell}$) depend on $k$, and (consequentially) the $k$-dependence of the norm of the solution operator is the same between any two spaces for which the solution operator is defined; i.e., the ${\cal H}_{0}\to{\cal H}_{0}$ norm, the ${\cal H}_{0}\to{\cal H}$ norm, and the ${\cal H}^{*}\to{\cal H}$ norm all have the same $k$-dependence – see Lemma 4.2 below.

### v2

# Preconditioning FEM discretisations of the high-frequency Helmholtz and Maxwell equations by either perturbing the coefficients or adding absorption

@@ Abstract
This paper investigates the following question: given a Galerkin matrix corresponding to a finite-element discretisation of either the Helmholtz or time-harmonic Maxwell equations with variable coefficients, suppose that the coefficients of the underlying PDE are perturbed; how good an approximate inverse (i.e., preconditioner) is the resulting Galerkin matrix to the original Galerkin matrix? An important special case is when the perturbation consists of adding absorption (in the spirit of “shifted Laplacian preconditioning”). The results of this paper improve the Helmholtz results in [18, 22] and extend these results to the time-harmonic Maxwell equations, confirming a conjecture in the recent preprint [29].
1 Introduction
1.1 Context and motivation
The design of good preconditioners for the linear systems arising from finite-element discretisations of high-frequency time-harmonic wave problems – such as the Helmholtz or time-harmonic Maxwell equations – is a longstanding open problem; see, e.g., the review articles [11, 14, 20, 19] and the references therein. This paper considers the theory of the following two, related, questions involving preconditioning these linear systems.
@@ 1.1 Context and motivation
The results of [18] were generalised in [22] to bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$, where $\mathsf{A}_{\ell}$, $\ell=1,2$, are the Galerkin matrices of $k^{-2}\nabla\cdot(A_{\ell}\nabla u_{\ell})+n_{\ell}u_{\ell}=f$. The motivation for studying this situation comes from uncertainty quantification (UQ): calculating quantities of interest for the solution of the Helmholtz equation $k^{-2}\nabla\cdot(A\nabla u)+nu=f$ with random coefficients requires the solution of many deterministic Helmholtz problems, each with different coefficients $A$ and $n$. Bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ then indicate to what extent a previously-calculated inverse of one of the Galerkin matrices can be used as a preconditioner for other Galerkin matrices; see [36, §4.6] and the recent work [39] for UQ algorithms using this idea.
@@
1.2 Informal statement of the main result
Set-up at the continuous level.
[…]
$k^{-2}\nabla\cdot(\mu_{\ell}^{-1}\nabla)-\epsilon_{\ell}$ (1.1)
or the time-harmonic Maxwell operator
$k^{-2}{\rm curl}\big(\mu^{-1}_{\ell}{\rm curl}\big)-\epsilon_{\ell}$ (1.2)
[…]
Specific Helmholtz and Maxwell problems to which the main result applies are described in Lemmas 2.2 and 2.4 below; we highlight that these include Helmholtz and Maxwell problems where the radiation condition is approximated by either a perfectly-matched layer (PML) or an impedance boundary condition.
Set-up at the discrete level.
Given a finite-dimensional subspace ${\cal H}_{N}\subset{\cal H}$ (with $N={\rm dim}({\cal H}_{N})$), let $\mathsf{A}_{\ell},\ell=1,2,$ be the Galerkin matrices associated to $a_{\ell}(\cdot,\cdot)$. Let the matrix $\mathsf{D}$ be such that
$\left\|v_{N}\right\|_{{\cal H}}^{2}=\left\|\mathbf{V}\right\|_{\mathsf{D}}^{2}\quad\text{ for }v_{N}=\sum_{j=1}^{N}V_{j}\phi_{j}\,\text{ with }\,V_{j}\in\mathbb{C}^{N};$ (1.3)
[…]
Informal statement of the main result (Theorem 4.2 below).
The main result gives sufficient conditions under which ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ (measured in either $\|\cdot\|_{\mathsf{D}}$ or the Euclidean norm $\|\cdot\|_{2}$) is controlled by $\|\mu_{1}-\mu_{2}\|_{L^{\infty}}$ and $\|\epsilon_{1}-\epsilon_{2}\|_{L^{\infty}}$ multiplied by the continuous inf-sup constant of $a_{1}(\cdot,\cdot)$.
In more detail, suppose that ${\cal H}_{N}$ is sufficiently large so that the discrete inf-sup constant of $a_{1}(\cdot,\cdot)$ is bounded by a constant (independent of $N$) multiplied by the continuous inf-sup constant of $a_{1}(\cdot,\cdot)$; denote the latter by $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}$ (see §3 below for an explanation of this notation). If
$\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{L^{\infty}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{L^{\infty}}\Big)C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$
then $\mathsf{A}_{2}^{-1}$ exists and
$\hskip 85.35826pt\leq 2\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{L^{\infty}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{L^{\infty}}\Big)C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (1.4)
Furthermore, if $\mu_{1}=\mu_{2}$, then there exists $C_{2}$ (independent of $N$ and $k$) such that
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq C_{2}\left\|\epsilon_{1}-\epsilon_{2}\right\|_{L^{\infty}}C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}},$ (1.5)
where $\|\cdot\|_{2}$ denotes the Euclidean norm on matrices (induced by the Euclidean norm on vectors).
1.3 Discussion
1.3.1 The main result specialised to preconditioning with absorption
In the set-up of §1.2, let $\mu_{2}=\mu_{1}$ and $\epsilon_{2}=(1+{\rm i}\alpha)\epsilon_{1}$ with $\alpha>0$ (i.e., $a_{2}(\cdot,\cdot)$ is formed by addition absorption to $a_{1}(\cdot,\cdot)$). If
$\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$
(i.e., the absorption is sufficiently small, depending on the inf-sup constant of $a_{1}(\cdot,\cdot)$), then $\mathsf{A}_{2}^{-1}$ exists,
and
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq C_{2}\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (1.7)
1.3.2 Recovering the Helmholtz results of [18] about preconditioning with absorption
[18] considers preconditioning the Helmholtz equation with absorption by letting $k^{2}\mapsto k^{2}+{\rm i}\varepsilon$. In the set-up above this corresponds to setting $\alpha:=\varepsilon/k^{2}$. By Remark 3.3 below, $\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\sim k$ when the Helmholtz problem is nontrapping. The right-hand side of (1.7) is then proportional to $\varepsilon/k$, which is the result proved in [18, Theorem 1.4] for the specific case of the Helmholtz interior impedance problem.
1.3.3 The condition that the discrete inf-sup constant is bounded by the continuous inf-sup constant
As stated in §1.2, the main result holds under the assumption that the dimension of the finite-dimensional approximation space is large enough so that the discrete inf-sup constant is bounded by a constant multiple of the continuous inf-sup constant.
This condition on the discrete inf-sup constant has been proven to hold for Helmholtz problems in the so-called asymptotic regime, i.e., when the finite-dimensional space is sufficiently large so that the Galerkin solution is quasi-optimal; see [32, Theorem 4.2] and the discussion in Remark 4.12 below. Recall that, for the $h$-version FEM applied to Helmholtz non-trapping problems, the asymptotic regime is when $(kh)^{p}k$ is sufficiently small [32, 6].
In §4.3 we prove a new result showing that the discrete inf-sup constant is bounded by a constant multiple of the continuous inf-sup constant in the preasymptotic regime when, for the $h$-version FEM applied to Helmholtz non-trapping problems, $(kh)^{2p}k$ is sufficiently small [17].
Given a Helmholtz or Maxwell problem of interest, the “recipe” to use the main result of this paper is therefore the following:
  1. 1.
    Show that the problem satisfies Assumption 2.1 below. Note that Lemmas 2.2 and 2.4 show that many commonly-studied Helmholtz and Maxwell problems satisfy this assumption.
  2. 2.
    Determine sufficient conditions for the discrete inf-sup constant to be bounded by a constant multiple of the continuous inf-sup constant – these conditions will involve regularity assumptions on the domain and coefficients, as well as conditions on the dimension of the approximation space. Section 4.3 below details these conditions (with references) for the $h$-FEM applied to the Helmholtz and Maxwell problems in Lemmas 2.2 and 2.4.
1.3.4 How the results of this paper improve the Helmholtz results of [18, 21]
Remark 4.7 below discusses in detail how the results of this paper improve the Helmholtz results of [18, 21]. We highlight here that the main way (aside from extending the results to the time-harmonic Maxwell equations) is in proving that bounds (1.4)-(1.7) hold if the discrete inf-sup constant is bounded by a constant multiple of the continuous inf-sup constant – the analyses in [18, 21] require more restrictive conditions on the finite-dimensional space.
1.3.5 Numerical experiments investigating the sharpness of the bounds (1.4), (1.5), (1.6), and (1.7)
[18, §5] and [22, §2.2.2] contain numerical experiments investigating the sharpness of the bounds (1.7) and (1.4)/(1.5) respectively. We highlight that these experiments consider piecewise-linear FEM discretisations in the preasymptotic regime, i.e., $(kh)^{2p}k$ with $p=1$; these FEM discretisations were not covered by the theory in [18, 22] (see the discussion in §1.3.3 and in Remark 4.7 below) but are covered by the results of the present paper.
1.3.6 The use of the bound (1.6) in the DD analyses of [27] and [29]
The Helmholtz two-level DD analysis of [27] assumes that the bound (1.6) holds for the Helmholtz interior impedance problem (see the assumptions of [27, Corollary 6.1]). Similarly, the Maxwell two-level DD analysis of [29] assumes that the bound (1.6) holds for the Maxwell interior impedance problem (see the discussion on [29, Page 2]). The main result of this paper, Theorem 4.2, coupled with results about the discrete inf-sup constant in §4.3 below then give sufficient conditions on the finite-element spaces for these assumptions to be satisfied.
Outline of the rest of the paper.
@@ 1.3.4 How the results of this paper improve the Helmholtz results of [18, 21]
§2 states the abstract assumptions under which the main results are proved and describes Helmholtz and Maxwell problems that are covered by these abstract assumptions. §3 recaps properties of the inf-sup constant of a sesquilinear form and its $k$-dependence for Helmholtz and Maxwell. §4 states the main results, and §5 proves them.
2 The class of Helmholtz and Maxwell problems considered
The main result is proved under the following abstract assumptions.
Assumption 2.1 (Abstract assumptions).
[…]
@@ 2 The class of Helmholtz and Maxwell problems considered
We make three immediate remarks.
@@
We now give examples of Helmholtz and Maxwell problems satisfying Assumption 2.1.
Lemma 2.2 (Helmholtz problems satisfying Assumption 2.1).
Let $\Omega$ be a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (scalar valued), ${\mathcal{D}}:=k^{-1}\nabla$, and
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}\nabla v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}.$ (2.4)
For $\ell=1,2,$ let $\mu_{\ell}^{-1}$ be bounded, symmetric, matrix functions on $\Omega$ with ${\rm essinf}_{\Omega}\Re\mu_{\ell}^{-1}>0$ (in the sense of quadratic forms) and $\epsilon_{\ell}$ be bounded, scalar-valued functions on $\Omega$.
If one of the following three points holds, then Assumption 2.1 holds.
[…]
and $b(\cdot,\cdot)=0$.
(ii) With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$, and $\Gamma_{\rm D}$ a (possibly empty) connected component of $\partial\Omega$ that is not equal to $\Gamma_{\rm imp}$,
${\cal H}:=\big\{v\in H^{1}(\Omega)\,:\,v=0\text{ on }\Gamma_{\rm D}\big\},$
@@
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$.
[…]
$b(u,v):=-{\rm i}k^{-1}\big({\rm DtN}_{k}u,v\big)_{L^{2}(\Gamma_{\rm imp})},$
where ${\rm DtN}_{k}$ is the exact Dirichlet-to-Neumann map for the Helmholtz equation in the exterior of $\{x:|x|\leq R\}$; see, e.g., [2, Equations 3.5 and 3.6] for explicit expressions in terms of Fourier series (in 2-d) or spherical harmonics (in 3-d) and Bessel and Hankel functions.
Remark 2.3 (The Helmholtz problems covered in Lemma 2.2).
The fact that ${\mathcal{D}}:=k^{-1}\nabla$ implies that the sesquilinear form (2.1) corresponds to the Helmholtz operator (1.1) Parts (i), (ii), and (iii) of Lemma 2.2 then cover three common ways of approximating the Sommerfeld radiation condition (i.e., the fact the scattered wave travels “outward” from the scatterer; see, e.g., [7, §3.2]).
Indeed, in Part (i), zero Dirichlet boundary conditions are imposed on at least one connected component of $\partial\Omega$, with then zero Neumann boundary conditions imposed on the rest. The requirements on $\mu_{\ell}$ and $\epsilon_{\ell}$ in Lemma 2.2 are then satisfied by the coefficients coming from a radial perfectly-matched layer (PML) by, e.g., [16, Lemma 2.3].
In Part (ii), the sesquilinear form corresponds to the Helmholtz operator (1.1) with the impedance boundary condition
@@
$u=0$ on $\Gamma_{\rm D}$, and $k^{-1}\partial_{n,\mu^{-1}}u=g_{\rm N}$ on $\partial\Omega\setminus(\Gamma_{\rm D}\cup\Gamma_{\rm imp})$ (where the data $g_{\rm imp}$ and $g_{\rm N}$ depend on the given right-hand side in the variational problem). Here $\partial_{n,\mu^{-1}}$ is the conormal derivative; recall that this is such that $\partial_{n,\mu^{-1}}u:=n\cdot\mu^{-1}\nabla u$ when $u\in H^{2}(\Omega)$, where $n$ is the outward-pointing unit normal vector to $\partial\Omega$ (see, e.g., [31, Lemmas 4.2 and 4.3]).
In Part (iii), the sesquilinear form corresponds to the Helmholtz operator (1.1) with the exact Dirichlet-to-Neumann map imposed on the boundary of a ball.
Proof of Lemma 2.2.
We need to check that (a) $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$, (b) $b(\cdot,\cdot)$ is continuous, and (c) the inequality (2.2) holds.
Regarding (a): this follows immediately from the definition of the norm (2.4) and the fact that ${\mathcal{D}}:=k^{-1}\nabla$.
Regarding (b): for Part (i), $b(\cdot,\cdot)=0$ and so is automatically continuous. For Part (ii), continuity of $b(\cdot,\cdot)$ follows from the standard multiplicative trace inequality (see, e.g., [24, Theorem 1.5.1.10, last formula on p. 41], [1, Theorem 1.6.6]). For Part (iii), continuity of $b(\cdot,\cdot)$ follows from, e.g., [32, Lemma 3.3, Part 1].
Regarding (c): for Part (i), the inequality (2.2) follows immediately since $b(\cdot,\cdot)=0$. For Part (ii), $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3) (with the term on $\Gamma_{\rm imp}$ playing no role because of the real part on the left-hand side of (2.3)), and thus (2.2) holds. For Part (iii), recall that $\Re(-{\rm i}{\rm DtN}_{k})\geq 0$ by, e.g., [2, Corollary 3.1] or [32, Lemma 3.3, Part 2], so that $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3). ∎
Lemma 2.4 (Maxwell problems satisfying Assumption 2.1).
Let $\Omega$ be a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (vector valued), and ${\mathcal{D}}:=k^{-1}{\rm curl}$.
For $\ell=1,2,$ let $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ be bounded, symmetric, matrix functions on $\Omega$ with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm essinf}_{\Omega}}}\Re\mu_{\ell}^{-1}>0$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm essinf}_{\Omega}}}\Re\epsilon_{\ell}>0$ (in the sense of quadratic forms).
If one of the following two points holds, then Assumption 2.1 holds.
@@
where $v_{T}$ is the tangential trace such that $v_{T}:=(n\times v)\times n$ for smooth $v$, where $n$ is the outward-pointing unit normal vector (see, e.g., [34, Equation 3.46 and Theorem 3.31]).
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}$ (2.5)
and $b(\cdot,\cdot)=0$.
(ii) With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$,
${\cal H}:=\big\{u\in H^{1}(\Omega)\,:\,\text{ $u_{T}=0$ on $\partial\Omega\setminus\Gamma_{\rm imp}$ and $u_{T}\in L^{2}(\Gamma_{\rm imp})$}\big\},$
$\left\|v\right\|^{2}_{{\cal H}}:=k^{-2}\left\|{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}+k^{-1}\left\|v_{T}\right\|^{2}_{L^{2}(\Gamma_{\rm imp})},$ (2.6)
@@
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$. In addition, $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are real-valued.
Remark 2.5 (The Maxwell problems covered in Lemma 2.2).
The fact that ${\mathcal{D}}:=k^{-1}{\rm curl}$ implies that the sesquilinear form (2.1) corresponds to the Maxwell operator and the sesquilinear form corresponds to the Maxwell operator (1.2) Parts (i) and (ii) of Lemma 2.2 then cover two common ways of approximating the radiation condition.
Indeed, the requirements on $\mu_{\ell}$ and $\epsilon_{\ell}$ in Part (i) are satisfied by the coefficients coming from a radial PML by, e.g., [4, Appendix A]. In Part (ii), the sesquilinear form corresponds to the Maxwell operator (1.2) with the impedance boundary condition
@@
Proof of Lemma 2.4.
As in the Helmholtz case, we need to check that (a) $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$, (b) $b(\cdot,\cdot)$ is continuous, and (c) the inequality (2.2) holds.
Regarding (a): this follows immediately from the definitions of the norm (2.5)/(2.6) and the fact that ${\mathcal{D}}:=k^{-1}{\rm curl}$.
Regarding (b): for Part (i), continuity of $b(\cdot,\cdot)$ is immediate (since it is zero). For Part (ii), continuity of $b(\cdot,\cdot)$ follows from the Cauchy–Schwarz inequality and the definition of the norm (2.6).
Regarding (c): for Part (i), the inequality (2.2) follows immediately since $b(\cdot,\cdot)=0$. For Part (ii), since $\mu^{-1}_{\ell}$ and $\epsilon_{\ell}$ are both real (by assumption), $\Im(a(v,v))$ consists only of $\Im(b(v,v))$, which is bounded below by $k^{-1}({\rm essinf}_{\Gamma_{\rm imp}}\theta)\|v\|^{2}_{L^{2}(\Gamma_{\rm imp})}$; the bound (2.2) then follows. ∎
@@
[…]